given y=4/6 and (-7,-7)
given answer -x-13 answer maybe off by 1 according to teacher he gave us the answer a head of time we have to find the parallel line of y=-x-13.i tried plugging in y-y1=m(x-x1)+b and plugging in the two sevens but i get lost after that

Answer by Alan3354(69443)

given y=4/6 and (-7,-7)
given answer -x-13 answer maybe off by 1 according to teacher he gave us the answer a head of time we have to find the parallel line of y=-x-13.i tried plugging in y-y1=m(x-x1)+b and plugging in the two sevens but i get lost after that
-----------
What is the problem?
What do you want to do?
You don't need to write a book, it's math, not Literature.